A data set lists full IQ scores for a random sample of subjects with low lead levels in their blood and another random sample of subjects with high lead levels in their blood. The statistics are summarized below. Use a 0.05 significance level to test the claim that the mean IQ score of people with low lead levels is higher than the mean IQ score of people with high lead levels.

Low Blood Lead Level: n = 77 x = 92.97403 s = 15.23239

High Blood Lead Level: n = 22 x = 87.45455 s = 9.246504

Assume that the two samples are independent simple random samples selected from normally distributed populations, and do not assume that the population standard deviation are equal. Test the given claim using the P-value method and critical value method.

Show the P Value, Test statistic, conclusion & confidence interval
Answer #1

To Test :-

H0 :- µ1 = µ2
H1 :- µ1 ≠ µ2

Test Statistic :-


t = 1.6806


Test Criteria :-
Reject null hypothesis if t > t(α, DF)


DF = 18
Critical value t(α, DF) = t( 0.05 , 18 ) = 1.734
t > t(α, DF) = 1.6806 < 1.734
Conclusion :- Fail to Reject Null Hypothesis


Decision based on P value
P - value = P ( t > 1.6806 ) = 0.0551
Reject null hypothesis if P value < α level of significance
P - value = 0.0551 > 0.05 ,hence we fail to reject null hypothesis
Conclusion :- Fail to Reject Null Hypothesis

Confidence interval :-

t(α/2, DF) = t(0.05 /2, 18 ) = 2.101

Lower Limit =
Lower Limit = -1.3806
Upper Limit =
Upper Limit = 12.4195
95% Confidence interval is ( -1.3806 , 12.4195 )

Since, interval contain the value 0, we fail to reject null hypothesis.